OpenStudy (anonymous):

Polygon ABCDE is dilated by a scale factor of 3 with the center of dilation at the origin to create polygon A′B′C′D′E′. If the endpoints of BC¯¯¯¯¯ are B(3, 5) and C(5, 10), what is the slope of B'C'¯¯¯¯¯¯¯ ?

ganeshie8 (ganeshie8):

dilation will not change slopes or angles so you can simply find the slope of BC

ganeshie8 (ganeshie8):

remember the slope formula ?

OpenStudy (anonymous):

yea

ganeshie8 (ganeshie8):

good, use it

ganeshie8 (ganeshie8):

B(3, 5) and C(5, 10)

OpenStudy (anonymous):

ok thank you i got (2,5)

ganeshie8 (ganeshie8):

slope is a single number, NOT a point..

OpenStudy (anonymous):

www.basic-mathematics.com/slope-calculator.html use this for your other slope questions

ganeshie8 (ganeshie8):

Oh wait, do you mean you get \(\large \dfrac{5}{2}\) for slope ??

OpenStudy (anonymous):

the formula is (X2-X1) (Y2-Y1)

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

slope formula is \[\large \dfrac{y_2-y_1}{x_2-x_1}\]

ganeshie8 (ganeshie8):

Notice there is a horizontal bar between (y2-y1) and (x2-x1) it refers to division

ganeshie8 (ganeshie8):

slope = (y2-y1) / (x2-x1)

OpenStudy (anonymous):

yeah

ganeshie8 (ganeshie8):

5/2 is right ! good job :)

OpenStudy (anonymous):

thx :D

ganeshie8 (ganeshie8):

yw!